# Avoiding Catastrophe in Online Learning by Asking for Help

### v3

@@ Appendix D There exist policy classes which are learnable in our setting but not in the standard online model
This section presents another algorithm with subconstant regret and sublinear queries, but under different assumptions. The most important aspect of our algorithm is that it can handle the class of thresholds on $[0,1]$, which is known to have infinite Littlestone dimension and thus be hard in the standard online learning model. (Example 21.4 in [34]).
Specifically, we assume a 1D state space and we allow the state sequence to be fully adversarial chosen. Instead of VC/Littlestone dimension, we consider the following notion of simplicity:
@@
Given a mentor policy $\pi^{m}$, partition the state space $S$ into intervals such that all states within each interval share the same mentor action. Let $\{S_{1},\dots,S_{k}\}$ be a partition that minimizes the number of intervals. We call each $S_{j}$ a segment. Let $f(\pi^{m})$ denote the number of segments in $\pi^{m}$.
Bounding the number of segments is similar conceptually to VC dimension in that it limits the ability of the policy class to realize arbitrary combinations of labels (i.e., mentor actions) on $\mathbf{s}$. For example, if $\Pi$ is the class of thresholds on $[0,1]$, every $\pi\in\Pi$ has at most two segments, and thus the positive result in this section will apply. This demonstrates the existence of policy classes which are learnable in our setting but not learnable in the standard online learning model, meaning that the two settings do not exactly coincide.
We prove the following result:
@@
[…]
@@
For any $c\in(1/2,1)$, Algorithm 4 with $g(T)=T^{c}$ makes $O(T^{c}(\operatorname*{diam}(\mathbf{s})+1))$ queries and satisfies
$\lim_{T\to\infty}\ \sup_{\mathbf{s}\in S^{T}}\ \sup_{\mu}\ \sup_{\pi^{m}:f(\pi^{m})\leq K}\ R_{T}=0$
@@ D.1 Intuition behind the algorithm
The algorithm maintains a set of buckets which partition the observed portion of the state space. Each bucket’s length determines the maximum loss in payoff we will allow from that subset of the state space. As long as the bucket contains a query from a prior time step, local generalization allows us to bound $\mu^{m}(s_{t})-\mu(s_{t},a_{t})$ based on the length of the bucket containing $s_{t}$. We always query if the bucket does not contain a prior query
@@
Regardless of the bucket length, the adversary can still place multiple segments in the same bucket $B$. A single query only tells us the optimal action for one of those segments, so we risk a payoff as bad as $\mu^{m}(s_{t})-O(\operatorname*{len}(B))$ whenever we choose not to query. We can endure a limited number of such payoffs, but if we never query again in that bucket, we may suffer $\Theta(T)$ such payoffs. Letting $\mu^{m}(s_{t})=1$ for simplicity, that would lead to $\prod_{t=1}^{T}\mu(s_{t},a_{t})\leq\big{(}1-\frac{1}{O(g(T))}\big{)}^{\Theta(T)}$, which converges to 0 (i.e., guaranteed catastrophe) when $g(T)\in o(T)$.
@@
2:     $S_{Q}\leftarrow\emptyset$triangleright Previously queried states
3:     $\pi\leftarrow\emptyset$triangleright Records $\pi^{m}(s)$ for each $s\in S_{Q}$
@@
6:         EvaluateState$(s_{t})$
7:function EvaluateState($s\in S$)
8:     if $s\not\in B$ for all $B\in\mathcal{B}$ then $\triangleright$ No bucket containing $s$: create a new bucket and try again
9:         $B\leftarrow\left[\frac{j-1}{g(T)},\frac{j}{g(T)}\right]$ for $j\in\mathbb{Z}$ such that $s\in B$
@@
14:         $B\leftarrow$ any bucket containing $s$
15:         if $S_{Q}\cap B=\emptyset$ then $\triangleright$ No queries in this bucket
16:              Query mentor and observe $\pi^{m}(s)$
17:              $\pi(s)\leftarrow\pi^{m}(s)$
18:              $S_{Q}\leftarrow S_{Q}\cup\{s\}$
@@
21:              Let $s^{\prime}\in S_{Q}\cap B$
22:              Take action $\pi(s^{\prime})$
@@
25:              $B=[x,y]$
26:              $(B_{1},B_{2})\leftarrow\Big{(}\left[x,\frac{x+y}{2}\right],\left[\frac{x+y}{2},y\right]\Big{)}$
27:              $(s_{B_{1}},s_{B_{2}})\leftarrow(0,0)$
@@ D.2 Notation for the proof
  • Let $V_{T}=\{t\in[T]:\mu(s_{t},a_{t})<\mu^{m}(s_{t})\}$ be the set of time steps with a suboptimal payoff.
@@
  • Let $\mathcal{B}_{V}=\{B:\exists t\in V_{T}\text{ s.t. }B_{t}=B\}$ be the set of buckets that ever produced a suboptimal payoff.
@@ D.4 Proof
Algorithm 4 performs at most $(\operatorname*{diam}(\mathbf{s})+4)g(T)$ queries.
@@
Since depth 0 buckets overlap only at their boundaries, and each depth 0 bucket has length $1/g(T)$, at most $g(T)\max_{t,t^{\prime}\in[T]}|s_{t}-s_{t^{\prime}}|=g(T)\operatorname*{diam}(\mathbf{s})$ depth 0 buckets are subsets of the interval $[\min_{t\in[T]}s_{t},\max_{t\in[T]}s_{t}]$. At most two depth 0 buckets are not subsets of that interval (one at each end), so the total number of depth 0 buckets is at most $g(T)\operatorname*{diam}(\mathbf{s})+2$.
We split a bucket $B$ when $n_{B}$ reaches $T/g(T)$, which creates two new buckets. Since each time step increments $n_{B}$ for a single bucket $B$, and there are a total of $T$ time steps, the total number of buckets created via splitting is at most
$\frac{2T}{T/g(T)}=2g(T)$
Therefore the total number of buckets ever in existence is $(\operatorname*{diam}(\mathbf{s})+2)g(T)+2\leq(\operatorname*{diam}(\mathbf{s})+4)g(T)$, so Algorithm 4 performs at most $(\operatorname*{diam}(\mathbf{s})+4)g(T)$ queries. ∎
@@
If we query the mentor at time $t$, $\mu(s_{t},a_{t})=\mu^{m}(s_{t})$. Thus assume we do not query the mentor at time $t$: then there exists $s^{\prime}\in B_{t}$ (as defined on line 21 of Algorithm 4) such that $a_{t}=\pi(s^{\prime})=\pi^{m}(s^{\prime})$. Since $s_{t}$ and $s^{\prime}$ are both in $B_{t}$, $|s_{t}-s^{\prime}|\leq\operatorname*{len}(B_{t})$. Then by the local generalization assumption,
$\mu(s_{t},a_{t})=\mu(s_{t},\pi^{m}(s^{\prime}))\geq\mu^{m}(s_{t})-L||s_{t}-s^{\prime}||\geq\mu^{m}(s_{t})-L\operatorname*{len}(B_{t})$
as required. ∎
@@
If two buckets $B$ and $B^{\prime}$ overlap by more than at a single point, one must be the descendant of the other. Thus by definition of $\mathcal{B}_{V}^{\prime}$, any two buckets in $\mathcal{B}_{V}^{\prime}$ overlap by at most a single point.
Now consider any $B\in\mathcal{B}_{V}^{\prime}$. By definition, there exists $t\in V_{T}$ such that $s_{t}\in B$. Then there exists $s^{\prime}\in B$ (as defined in Algorithm 4) such that $a_{t}=\pi(s^{\prime})=\pi^{m}(s^{\prime})$. Since $t\in V_{T}$, we have $\pi^{m}(s_{t})\neq a_{t}=\pi^{m}(s^{\prime})$. Thus $s_{t}$ and $s^{\prime}$ are in different segments; let $S_{B}$ be the segment containing $\min(s_{t},s^{\prime})$. Then $\min(s_{t},s^{\prime})\leq\max(S_{B})<\max(s_{t},s^{\prime})\leq\max(B)$.
Consider any $B,B^{\prime}\in\mathcal{B}_{V}^{\prime}$ with $B\neq B^{\prime}$ and WLOG assume $\max(B)<\max(B^{\prime})$. Then $\max(S_{B})<\max(B)\leq\max(S_{B^{\prime}})$, so $S_{B}\neq S_{B^{\prime}}$. Thus the number of segments in $\pi^{m}$ is at least $\mathcal{B}_{V}^{\prime}$. Since the number of segments in $\pi^{m}$ is at most $K$, we must have $|\mathcal{B}_{V}^{\prime}|\leq K$. ∎
@@
For every $t\in V_{T}$, we have $B_{t}=B$ for some $B\in\mathcal{B}_{V}$, so
$\sum_{t\in V_{T}}\operatorname*{len}(B_{t})=\sum_{B\in\mathcal{B}_{V}}\ \sum_{t\in V_{T}:B=B_{t}}\operatorname*{len}(B_{t})$
@@
$\sum_{t\in V_{T}}$ $\ \operatorname*{len}(B_{t})\leq\sum_{B^{\prime}\in\mathcal{B}_{V}^{\prime}}\ \sum_{B\in\mathcal{A}(B^{\prime})}\ \sum_{t\in V_{T}:B=B_{t}}\operatorname*{len}(B_{t})$
$=$ $\ \sum_{B^{\prime}\in\mathcal{B}_{V}^{\prime}}\ \sum_{B\in\mathcal{A}(B^{\prime})}\ |\{t\in V_{T}:B=B_{t}\}|\cdot\operatorname*{len}(B_{t})$
@@
$\sum_{B\in\mathcal{A}(B^{\prime})}$ |$\ \frac{|\{t\in V_{T}:B=B_{t}\}|}{g(T)2^{d(B)}}\leq\frac{T}{g(T)^{2}}\sum_{B\in\mathcal{A}(B^{\prime})}\ \frac{1}{2^{d(B)}}$
@@
$\sum_{t\in V_{T}}\operatorname*{len}(B_{t})\leq\sum_{B^{\prime}\in\mathcal{B}_{V}^{\prime}}\frac{2T}{g(T)^{2}}=\frac{2T|\mathcal{B}_{V}^{\prime}|}{g(T)^{2}}\leq\frac{2KT}{g(T)^{2}}$
@@
For $t\not\in V_{T}$ we have $\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))=\mu^{m}(s_{t})$ by definition, and Lemma D.2 implies that $\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))\geq L\operatorname*{len}(B_{t})$ for all $t\in[T]$. Thus
$\sum_{t=1}^{T}\Big{(}\mu^{m}(s_{t})-\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))\Big{)}\leq\sum_{t\in V_{T}}\Big{(}\mu^{m}(s_{t})-\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))\Big{)}\leq L\sum_{t\in V_{T}}\operatorname*{len}(B_{t})$
@@
Under the conditions of Theorem D.2, Algorithm 4 satisfies $R_{T}\leq\frac{2LKT}{g(T)^{2}}$.
@@
Let $x_{t}=\mu^{m}(s_{t})$ and $y_{t}=\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))$ for all $t\in[T]$. Then by Lemma B.3,
$\prod_{t=1}^{T}\mu^{m}(s_{t})-\prod_{t=1}^{T}\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))\leq$ $\ \sum_{t=1}^{T}\Big{(}\mu^{m}(s_{t})-\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))\Big{)}$
Since $\mu(s_{t},a_{t})\geq\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))$ for all $t\in[T]$, we have
$R_{T}=\prod_{t=1}^{T}\mu^{m}(s_{t})-\prod_{t=1}^{T}\mu(s_{t},a_{t})\leq\sum_{t=1}^{T}\Big{(}\mu^{m}(s_{t})-\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))\Big{)}$
Applying Lemma D.5 completes the proof. ∎

### v4

@@ Appendix D There exist policy classes which are learnable in our setting but not in the standard online model
This section presents another algorithm with subconstant regret and sublinear queries, but under different assumptions. The primary takeaway here is that our algorithm can handle the class of thresholds on $[0,1]$, which is known to have infinite Littlestone dimension and thus be hard in the standard online learning model. (Example 21.4 in Shalev-Shwartz & Ben-David (2014)).
Specifically, we assume a 1D input space and we allow the input sequence to be fully adversarial chosen. Instead of VC/Littlestone dimension, we consider the following notion of simplicity:
@@
Given a mentor policy $\pi^{m}$, partition the input space $\mathcal{X}$ into intervals such that all inputs within each interval share the same mentor action. Let $\{X_{1},\dots,X_{k}\}$ be a partition that minimizes the number of intervals. We call each $X_{j}$ a segment. Let $S(\pi^{m})$ denote the number of segments in $\pi^{m}$.
Bounding the number of segments is similar conceptually to VC dimension in that it limits the ability of the policy class to realize arbitrary combinations of labels (i.e., mentor actionx) on $\boldsymbol{x}$. For example, if $\Pi$ is the class of thresholds on $[0,1]$, every $\pi\in\Pi$ has at most two segments, and thus the positive result in this section will apply. This demonstrates the existence of policy classes which are learnable in our setting but not learnable in the standard online learning model, meaning that the two settings do not exactly coincide.
We prove the following regret bound. Like our other positive results, this bound applies to both multiplicative and additive regret.
@@
For any $\boldsymbol{x}\in\mathcal{X}^{T}$, any $\pi^{m}$ with $S(\pi^{m})\leq K$, and any function $g:\mathbb{N}\to\mathbb{N}$, Algorithm 4 makes at most $(\operatorname*{diam}(\boldsymbol{x})+4)g(T)$ queries and satisfies
$R_{T}\leq$ $\ \frac{2LKT}{g(T)^{2}}$
$\sum_{t=1}^{T}\mu_{t}^{m}(x_{t})-\sum_{t=1}^{T}\mu_{t}(x_{t},y_{t})\leq$ $\ \frac{2LKT}{g(T)^{2}}$
@@
For any $c\in(1/2,1)$, Algorithm 4 with $g(T)=T^{c}$ makes $O(T^{c}(\operatorname*{diam}(\boldsymbol{x})+1))$ queries and satisfies
$R_{T}\leq$ $\ 2LKT^{1-2c}$
$\sum_{t=1}^{T}\mu_{t}^{m}(x_{t})-\sum_{t=1}^{T}\mu_{t}(x_{t},y_{t})\leq$ $\ 2LKT^{1-2c}$
@@ D.1 Intuition behind the algorithm
The algorithm maintains a set of buckets which partition the observed portion of the input space. Each bucket’s length determines the maximum loss in payoff we will allow from that subset of the input space. As long as the bucket contains a query from a prior time step, local generalization allows us to bound $\mu_{t}^{m}(x_{t})-\mu_{t}(x_{t},y_{t})$ based on the length of the bucket containing $x_{t}$. We always query if the bucket does not contain a prior query
@@
Regardless of the bucket length, the adversary can still place multiple segments in the same bucket $B$. A single query only tells us the optimal action for one of those segments, so we risk a payoff as bad as $\mu_{t}^{m}(x_{t})-O(\operatorname*{len}(B))$ whenever we choose not to query. We can endure a limited number of such payoffs, but if we never query again in that bucket, we may suffer $\Theta(T)$ such payoffs. Letting $\mu_{t}^{m}(x_{t})=1$ for simplicity, that would lead to $\prod_{t=1}^{T}\mu_{t}(x_{t},y_{t})\leq\big{(}1-\frac{1}{O(g(T))}\big{)}^{\Theta(T)}$, which converges to 0 (i.e., guaranteed catastrophe) when $g(T)\in o(T)$.
@@
2:     $X_{Q}\leftarrow\emptyset$triangleright Previously queried inputs
3:     $\pi\leftarrow\emptyset$triangleright Records $\pi^{m}(x)$ for each $x\in X_{Q}$
@@
6:         EvaluateInput$(x_{t})$
7:function EvaluateInput($x\in\mathcal{X}$)
8:     if $s\not\in B$ for all $B\in\mathcal{B}$ then $\triangleright$ No bucket containing $x$: create a new bucket and try again
9:         $B\leftarrow\left[\frac{j-1}{g(T)},\frac{j}{g(T)}\right]$ for $j\in\mathbb{Z}$ such that $x\in B$
@@
14:         $B\leftarrow$ any bucket containing $x$
15:         if $X_{Q}\cap B=\emptyset$ then $\triangleright$ No queries in this bucket
16:              Query mentor and observe $\pi^{m}(x)$
17:              $\pi(x)\leftarrow\pi^{m}(x)$
18:              $X_{Q}\leftarrow X_{Q}\cup\{x\}$
@@
21:              Let $x^{\prime}\in X_{Q}\cap B$
22:              Take action $\pi(x^{\prime})$
@@
25:              $B=[a,b]$
26:              $(B_{1},B_{2})\leftarrow\Big{(}\left[a,\frac{a+b}{2}\right],\left[\frac{a+b}{2},b\right]\Big{)}$
27:              $(x_{B_{1}},x_{B_{2}})\leftarrow(0,0)$
@@ D.2 Notation for the proof
  1. 1.
    Let $M_{T}=\{t\in[T]:\mu_{t}(x_{t},y_{t})<\mu_{t}^{m}(x_{t})\}$ be the set of time steps with a suboptimal payoff.
  2. 2.
@@
  5. 5.
    Let $\mathcal{B}_{V}=\{B:\exists t\in M_{T}\text{ s.t. }B_{t}=B\}$ be the set of buckets that ever produced a suboptimal payoff.
  6. 6.
@@ D.4 Proof
Algorithm 4 performs at most $(\operatorname*{diam}(\boldsymbol{x})+4)g(T)$ queries.
@@
Since depth 0 buckets overlap only at their boundaries, and each depth 0 bucket has length $1/g(T)$, at most $g(T)\max_{t,t^{\prime}\in[T]}|x_{t}-x_{t^{\prime}}|=g(T)\operatorname*{diam}(\boldsymbol{x})$ depth 0 buckets are subsets of the interval $[\min_{t\in[T]}x_{t},\max_{t\in[T]}x_{t}]$. At most two depth 0 buckets are not subsets of that interval (one at each end), so the total number of depth 0 buckets is at most $g(T)\operatorname*{diam}(\boldsymbol{x})+2$.
[…]
@@
If we query the mentor at time $t$, $\mu_{t}(x_{t},y_{t})=\mu_{t}^{m}(x_{t})$. Thus assume we do not query the mentor at time $t$: then there exists $x^{\prime}\in B_{t}$ (as defined on line 21 of Algorithm 4) such that $y_{t}=\pi(x^{\prime})=\pi^{m}(x^{\prime})$. Since $x_{t}$ and $x^{\prime}$ are both in $B_{t}$, $|x_{t}-x^{\prime}|\leq\operatorname*{len}(B_{t})$. Then by the local generalization assumption, $\mu_{t}(x_{t},y_{t})=\mu_{t}(x_{t},\pi^{m}(x^{\prime}))\geq\mu_{t}^{m}(x_{t})-L||x_{t}-x^{\prime}||\geq\mu_{t}^{m}(x_{t})-L\operatorname*{len}(B_{t})$. ∎
@@
Now consider any $B\in\mathcal{B}_{V}^{\prime}$. By definition of $\mathcal{B}_{V}^{\prime}$, there exists $t\in M_{T}$ such that $x_{t}\in B$. Then there exists $x^{\prime}\in B$ (as defined in Algorithm 4) such that $y_{t}=\pi(x^{\prime})=\pi^{m}(x^{\prime})$. Since $t\in M_{T}$, we have $\pi^{m}(x_{t})\neq y_{t}=\pi^{m}(x^{\prime})$. Thus $x_{t}$ and $x^{\prime}$ are in different segments, but are both in $B$. Therefore any $B\in\mathcal{B}_{V}^{\prime}$ must intersect at least two segments. Since $B$ is an interval, if it intersects two segments, it must intersect two adjacent segments $X_{j}$ and $X_{j+1}$. Furthermore, $B$ must contain an open neighborhood centered on the boundary between $X_{j}$ and $X_{j+1}$.
Now consider some $B^{\prime}\in\mathcal{B}_{V}^{\prime}$ with $B\neq B^{\prime}$. We $|B\cap B^{\prime}|\leq 1$: otherwise one must be the descendant of the other, which contradicts the definition of $\mathcal{B}_{V}^{\prime}$. Suppose $B^{\prime}$ also intersects both $X_{j}$ and $X_{j+1}$: since $B^{\prime}$ is also an interval, $B^{\prime}$ must also contain an open neighborhood centered on the boundary between those two segments. But then $|B\cap B^{\prime}|>1$, which is a contradiction.
Therefore any pair of adjacent segments $X_{j}$ and $X_{j+1}$, there is at most one bucket in $\mathcal{B}_{V}^{\prime}$ which contains an open neighborhood around their boundary. Since there are at most $K-1$ pairs of adjacent segments, we have $|\mathcal{B}_{V}^{\prime}|\leq K-1\leq K$. ∎
@@
For every $t\in M_{T}$, we have $B_{t}=B$ for some $B\in\mathcal{B}_{V}$, so
$\sum_{t\in M_{T}}\operatorname*{len}(B_{t})=\sum_{B\in\mathcal{B}_{V}}\ \sum_{t\in M_{T}:B=B_{t}}\operatorname*{len}(B_{t})$
@@
$\sum_{t\in M_{T}}$ $\ \operatorname*{len}(B_{t})\leq\sum_{B^{\prime}\in\mathcal{B}_{V}^{\prime}}\ \sum_{B\in\mathcal{A}(B^{\prime})}\ \sum_{t\in M_{T}:B=B_{t}}\operatorname*{len}(B_{t})$
$=$ $\ \sum_{B^{\prime}\in\mathcal{B}_{V}^{\prime}}\ \sum_{B\in\mathcal{A}(B^{\prime})}\ |\{t\in M_{T}:B=B_{t}\}|\cdot\operatorname*{len}(B_{t})$
@@
$\sum_{B\in\mathcal{A}(B^{\prime})}$ |$\ \frac{|\{t\in M_{T}:B=B_{t}\}|}{g(T)2^{d(B)}}\leq\frac{T}{g(T)^{2}}\sum_{B\in\mathcal{A}(B^{\prime})}\ \frac{1}{2^{d(B)}}$
@@
$\sum_{t\in M_{T}}\operatorname*{len}(B_{t})\leq\sum_{B^{\prime}\in\mathcal{B}_{V}^{\prime}}\frac{2T}{g(T)^{2}}=\frac{2T|\mathcal{B}_{V}^{\prime}|}{g(T)^{2}}\leq\frac{2KT}{g(T)^{2}}$
@@
For $t\not\in M_{T}$ we have $\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))=\mu_{t}^{m}(x_{t})$ by definition, and Lemma D.2 implies that $\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))\geq L\operatorname*{len}(B_{t})$ for all $t\in[T]$. Thus
$\sum_{t=1}^{T}\Big{(}\mu_{t}^{m}(x_{t})-\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))\Big{)}\leq$ $\ \sum_{t\in M_{T}}\Big{(}\mu_{t}^{m}(x_{t})-\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))\Big{)}$
$\leq$ $\ L\sum_{t\in M_{T}}\operatorname*{len}(B_{t})$
@@
Under the conditions of Theorem D.2, Algorithm 4 satisfies
$R_{T}\leq$ $\ \frac{2LKT}{g(T)^{2}}$
$\sum_{t=1}^{T}\mu_{t}^{m}(x_{t})-\sum_{t=1}^{T}\mu_{t}(x_{t},y_{t})\leq$ $\ \frac{2LKT}{g(T)^{2}}$
@@
Let $a_{t}=\mu_{t}^{m}(x_{t})$ and $b_{t}=\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))$ for all $t\in[T]$. Then by Lemma B.3,
$\prod_{t=1}^{T}\mu_{t}^{m}(x_{t})-\prod_{t=1}^{T}\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))\leq$ $\ \sum_{t=1}^{T}\Big{(}\mu_{t}^{m}(x_{t})-\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))\Big{)}$
Since $\mu_{t}(x_{t},y_{t})\geq\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))$ for all $t\in[T]$, we have
$R_{T}=\prod_{t=1}^{T}\mu_{t}^{m}(x_{t})-\prod_{t=1}^{T}\mu_{t}(x_{t},y_{t})\leq\sum_{t=1}^{T}\Big{(}\mu_{t}^{m}(x_{t})-\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))\Big{)}$
Similarly,
$\sum_{t=1}^{T}\mu_{t}^{m}(x_{t})-\sum_{t=1}^{T}\mu_{t}(x_{t},y_{t})\leq\sum_{t=1}^{T}\Big{(}\mu_{t}^{m}(x_{t})-\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))$
Applying Lemma D.5 to the above two inequalities completes the proof. ∎